OpenStudy (anonymous):

Help me will fan and medal. lx+3l=4

OpenStudy (anonymous):

solve two equations both take one step \[x+3=4\\ x+3=-4\]

OpenStudy (uri):

x+3=4 x+3=-4 solve these two

OpenStudy (uri):

k @satellite73 you win!